OpenStudy (anonymous):

the magnetic field at the point of intersection of the diagonals of a square loop of side L carrying a current i is?

OpenStudy (anonymous):

\[a)\mu i/\pi L\] \[b)2 \mu I/\pi L\] c)\[\sqrt{2} \mu I/\pi L\] d)\[2\sqrt{2} \mu I/\pi L\]

OpenStudy (anonymous):

could you maybe draw what you mean? :) i might be able to help

OpenStudy (ghazi):

it's D.....see the figure attached ...